Draw the Lewis structure for ICl5 and answer the following questions.



How many valence electrons are present in this compound?

How many bonding electrons are present in this compound?

How many lone pair (non-bonding) electrons are present in this compound?
Answer #1

image from custom entry tool

Add 6 dots around each Cl. 2 dots/1 lone pair on the Iodine for a total of 42electrons

How many valence electrons are present in this compound

7 from Iodine. 7X5=35 from Cl. Total 35+7=42

How many bonding electrons are present in this compound?
5

How many lone pair (non-bonding) electrons are present in this compound

1
Answer #3

The VESPR shape of this molecule is octahedral, with one Cl at the bottom and 4 evenly spaced around the I at the middle. I can accept more than an octet in electrons (12), so it has a lone pair up at the top.

Each Cl atom should have 3 pairs of electrons around it, since Cl DOES follow the octet rule. 3 * 2 = 6, + 2 shared electrons that come from the bond with I

Iodine has 7 val e-, 5e - making bonds with 5 Cl and 2 e- forms a lone pair.

6 electron domains - 6 bonds - octahedral
6 electron domains - 5 bonds and 1 lone pair (located either below or above a square plane) - square pyramidal
6 electron domains - 4 bonds and 2 lone pair (located below and above a square plane) - square planar.

ICl5 : First draw the Lewis dot structure:
ICl5 Lewis Structure
Electron geometry: octahedral. Hybridization: sp3d2.